OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. B(0, 0), C(4, -4)

OpenStudy (anonymous):

x+y=0

OpenStudy (anonymous):

but that cant be the answer

OpenStudy (anonymous):

Never mind that is the correct answer but could i get how u got that

OpenStudy (anonymous):

calculate the slope:y-y0/x-x0 put in y=mx+c where m is slope.and if my answer is not right maybe its because of c i assumed it to be 0.

OpenStudy (texaschic101):

m = (y2 - y1)/(x2 - x1)....set 1 (0,0)...set 2 (4,-4) m = (-4 - 0)/ (4 - 0) m = -4/4 = -1 y = mx + b ....using points (4,-4) -4 = -1(4) + b -4 = -4 + b -4 + 4 = b 0 = b y = -1x + 0 ax + by = c -x + y = 0 You might better get a second opinion on this